OpenStudy (anonymous):

simplify using only positive exponents x^-3+y^-3 over x^-2+y^-2

OpenStudy (anonymous):

x^-3 = 1/x³ Can you do similarly for the rest?

OpenStudy (anonymous):

what do you do after you've changed it to 1/x^3 +1/y^3 over 1/x^2+1/y^2

OpenStudy (anonymous):

if that is even right

OpenStudy (kropot72):

\[((2+(x/y)^{3}+(y/x)^{3})\div((x+y+x(x/y)^{2}+y(y/x)^{2}\] This was obtained by multiplying numerator an denominator by: \[x ^{3}+y ^{3}\]

OpenStudy (anonymous):

Can you make common denominator for them: 1/x^3 +1/y^3

OpenStudy (anonymous):

also common denom. 1/x^2 + 1/y^2

OpenStudy (anonymous):

There might be cleverer ways to do this, but I usually just go with the grind it out method....\[\frac{x^{-3}+y^{-3}} { x^{-2}+y^{-2} }=\frac{\frac{1}{x^3}+\frac{1}{y^3}}{\frac{1}{x^2}+\frac{1}{y^2}}=\frac{\frac{y^3+x^3}{x^3y^3}}{\frac{y^2+x^2}{x^2y^2}}=\frac{y^3+x^3}{xy(y^2+x^2)}\]